Solve by the elimination method
0.11x+0.09y=-0.7
0.3x-0.6y=5.7
What is the solution of the system?

Answer by checkley71(8403)

multiply eq #1 by 30 & the eq #2 by -11 we get
3.3x+2.7y=-21
-3.3x+6.6y=-62.7
--------------------
9.3y=-83.7 or y=-83.7/9.3 or y=-9 then .11x+.09*-9=-.7 or .11x-.81=-.7 or
.11x=.11 or x=1
proof .3*1-.6*-9=5.7 or .3+5.4=5.7 or 5.7=5.7